OpenStudy (anonymous):

So today on a test I was given this question thermodynamics: So I was given a formula, I think it was C4H8 +12O2 -> 4CO2 + 8H2O Or something along those lines Then I was given all the heat of formations. except for O2 and C4H8, so I calculated what the Hf was since I had the Enthalpy. Then an odd question stuck me: It gave me the mass of O2 then had me solve how much heat is produced: what could it be? Since I was not given the Specific heat of O2, I couldn't calculate the heat from that. So I converted it to moles and found the molar ratio of the other substances...

OpenStudy (anonymous):

I also had the bond enthalpies for each of the molecules... What should i have done?

OpenStudy (anonymous):

You don't need the information for the oxygen, because it is an element in its standard state. \[\Delta \text{H}^{\circ}_f = [\sum (\text{mol})(\Delta \text{H}^{\circ}_f(\text{products})] - [\sum (\text{mol})(\Delta \text{H}^{\circ}_f(\text{reactants})]\] If I understand correctly, based on your description, you should just be able to do this without requiring bond enthalpies.

OpenStudy (anonymous):

So... I wasn't suppose to consider the amount of oxygen at all? @ghuczek

OpenStudy (anonymous):

pellet, no one else has anything to add??

OpenStudy (aaronq):

the equation posted above is what you should've used to solve the question, though for oxygen \(H^o_f=0\) because it's in natural state (or whatever you wanna call it) no energy is needed to obtain it. For the part when they give you the mass of oxygen, you assume that you have an equivalent amount of moles of whatever you're burning (can't make heat with just oxygen). So you would find the moles, divide them by it's coefficient and multiply it by the enthalpy change of the reaction: \(\Delta H_{reaction}=\dfrac{n_{O_2}*\Delta H^o_{reaction}}{\nu_{O_2}}\)

OpenStudy (anonymous):

Thank you ... is there any additional explanation for it? So if they give me the general enthalpy, all the ratio must match up, so just giving me one will give me the ratio for the whole thing... thanks.

OpenStudy (aaronq):

yeah, i mean, unless they give you both reactants, then you have work use the moles of the limiting reactant.

OpenStudy (aaronq):

using*

OpenStudy (anonymous):

alright, thank you!